OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 8/x^2 + 4

OpenStudy (anonymous):

g(x) = x^2 and f(x) = ?

OpenStudy (anonymous):

f(x) = 8 or 12? still confused

OpenStudy (anonymous):

Substitute u=x^2 in your equation above. y=f(g(x))=f(u)=...

OpenStudy (anonymous):

So i would multiply 8 by x^2? BTW thank you for the help!

OpenStudy (anonymous):

No, it looks like in your equation above you have 8/x^2+4=8/u+4

OpenStudy (anonymous):

I just gave you the answer. Whoops!

OpenStudy (anonymous):

That would be the final answer? hm i was thinking totally something different...

OpenStudy (anonymous):

ohhh i gotcha thank you so much